Figure 1. Parameter space in the minimal supersymmetric model. Only two of the five dimensions (μ and M_{2}) are displayed. Panel (a) shows the starting grid of parameters choices, and panel (c) shows the models left after eliminating models which violate any of several accelerator constraints (from [22]).

Figure 2. Scatter plot of relic neutralino density vs. neutralino mass for the set of supersymmetric models discussed in the text. Laboratory constraints from LEP measurements and Br(b → sγ) are enforced. Models between the lines drawn at Ω_{χ}h^{2} = 0.025 and Ω_{χ}h^{2} = 1 are compatible with neutralino dark matter (from [22]).

Figure 3. Predicted rate in a ^{73}Ge cryogenic detector vs neutralino mass for the allowed dark-matter models above (from [22]).

Figure 4. Measured gamma-ray background in an underground high-purity germanium ionization detector (data acquired by the UCB/UCSB/LBL experiment at Oroville (from [26, 27]). Various gamma-ray lines are identified, as is the end point of the broad tritium spectrum. The rapid rise at low Q is the electronic noise (from [22]).

Figure 5. Theoretical differential event rate for WIMPS of various masses hitting a germanium target. WIMP masses are labeled in GeV. An arbitrary cross section of σ_{0} = 4 × 10^{−36}cm^{2} was chosen with standard values for the other parameters. Note the rate axis scale is 100 times smaller than in Fig. 4, and the cross section chosen is very high for neutralinos (from [22]).

Figure 6. Indirect-detection rate vs neutralino mass. The sum of the rates for upward muons from both the Sun and Earth is shown. Currently planned experiments will be sensitive in the 10^{−2}m^{−2}yr^{−1} to 10^{−4}m^{−2}yr^{−1} range (from [22]).

Figure 7. The three observed stellar lightcurves that we interpret as gravitational microlensing events are each shown in relative flux units (red and blue) vs time in days. The solid lines are fits to the theoretical microlensing shape (from [37]).

Figure 8. Example lightcurves from first year bulge data. Four of the 43 microlensing events are shown (from [46, 47]).

Figure 9. The lower panel shows the number of expected events predicted from the standard model halo with a delta function mass distribution. Given three observed events, points above the line drawn at N_{exp} = 7.7 are excluded at the 95% CL. The upper panel shows the 95% CL limit on the halo mass in MACHOs within 50 kpc of the galactic center for the model. Points above the curve are excluded at 95% CL while the line at 4.1 × 10^{11}M_{⊙} shows the total mass in this model within 50 kpc (from [37]).